What is the name of the following inorganic complexes

a. [Pt(NH3)3C13] b. [Co(NH3)2(H20)2Cl2] c. [Re(dien)Br2Cl] d. [Co(EDTA)] e. [Cr(phen) (NH3)2C12]*
Answer #1

@ Let me be the oxudation number of Pt. + x + 3X0 + 3 X (-1) = +1 + 2-3 = +1 >121 = +41 buamminebrichioridoplatinum (IV) het③ Let a be the oxidation number of Crotone + x + 140+240 +2 X(-1) = +1 bodo 1 -2 = t1 => 1 = +37 be con todos diamminedichlor